OpenStudy (xbeenchantedx):

Rewrite the quadratic equation in vertex form, f(x)=a(x-h)^2+k, by completing the square. j(x)=2x^2-11x-18

OpenStudy (anonymous):

do u rmember how to do the completing the square?

OpenStudy (xbeenchantedx):

No, I'm a little rusty on that.

OpenStudy (anonymous):

if i just give u that. can u understand?

OpenStudy (anonymous):

but. at first. u must divide the whole eq with the 'a' in ax2

OpenStudy (anonymous):

then,the first step u will got this. \[2(x ^{2} - \frac{ 11 }{ 2 }x -\frac{ 18 }{ 2 })\]

OpenStudy (anonymous):

next, \[2(x ^{2} - \frac{ 11 }{ 2 }x = 9)\]

OpenStudy (anonymous):

then, u may continue with the help of attachment that i had gave u early. u try it first n tell me the ans

OpenStudy (xbeenchantedx):

i have absolutely no idea how I even got (x-11/2)^2=47/4 but that's what I got.

OpenStudy (anonymous):

that is called completing the square formula.. u want or not. the formula is like that. :D